# Large deviations for the smallest eigenvalue of a deformed GOE with an outlier

### v1

@@ 2.3. Outliers of $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$
$\frac{y_{0}}{\Lambda_{1}-\Lambda}+\sum_{k=1}^{p}\frac{y_{k}}{\Lambda_{1}-\eta_{k}}=0.$
Item (2) is clear. Moreover, the continuity of $\Phi(\Lambda,Y)$ is clear from (2.12). Indeed, if $\varepsilon>0$ is such that $Y\subset[\varepsilon,1-\varepsilon]\times[0,1]^{p}\cap\Delta^{p+1}$, we see from (2.12) that
$\varepsilon(\eta_{1}-\Lambda_{1})\leq\Lambda_{1}-\Lambda\leq\varepsilon^{-1}(\eta_{1}-\Lambda_{1})$
so that $\Lambda_{1}$ is at distance of order $\varepsilon$ of $\Lambda$ and $\eta_{1}$. Then, if $Y,Y^{\prime}\in[\varepsilon,1-\varepsilon]\times[0,1]^{p}\cap\Delta^{p+1}$, we easily deduce that
$|\Phi(\Lambda,Y)-\Phi(\Lambda,Y^{\prime})|\leq C(\varepsilon)\sum_{i=1}^{p}|y_{i}-y_{i}^{\prime}|\,.$
When $y_{0}$ goes to zero or one, it is clear that $\Phi(\Lambda,Y)$ was extended by continuity, which in turn completes the proof of the continuity of $\Phi(\Lambda,.)$. ∎
We next study the typical behavior of the smallest eigenvalue of $p_{v_{1}}^{\perp}(G_{N}+B_{N})p_{v_{1}}^{\perp}$ in $v_{1}^{\perp}$, when $v_{1}$ is as in Lemma 2.4. It is a straightforward application of Lemma 2.4 and Theorem 1.1 as this matrix is the sum of a GOE matrix and the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ with spectrum asymptotically distributed according to $\nu$ and outlier $\Phi(Y,\Lambda)$ as in Lemma 2.4.
@@ 2.4. A functional inequality for the large deviation upper bound
Let $B_{N}$ be a diagonal matrix with eigenvalues $\Lambda$ with multiplicity $1$, $\eta_{1},\ldots,\eta_{p}$ of multiplicities $M_{1},\ldots,M_{p}$. Let $\mathsf{M}=(M_{1},\ldots,M_{p})\in\mathcal{B}_{\delta}^{N}(\alpha)$. For each $\Lambda\leq\ell_{\nu}$, let us denote $\mathbb{P}_{N,\Lambda,\mathsf{M}}$ the law of $X_{N}=G_{N}+B_{N}$ and
@@
$F_{\nu,t}^{+}(\Lambda,\lambda):=-\limsup_{\Lambda^{\prime}\to\Lambda}\limsup_{\varepsilon\to 0}\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda,\mathsf{M}}(\lambda,\varepsilon).$ (2.14)
@@
$g:x\in\mathbb{R}\mapsto\lim_{\varepsilon\to 0}\sup_{B(x,\varepsilon)}f$
@@
We consider the map $L_{\nu,t}^{\Lambda}:(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\to\mathbb{R}$, given for all $\lambda\in(-\infty,\ell_{\nu,t}]$ and $Y\in\Delta^{p+1}$ by
Recall the map $I_{\nu}:Y\in\Delta^{p+1}\mapsto-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\log(y_{k}/\alpha_{k})$ from Lemma 2.3. We set
$J_{\nu,t}^{\Lambda}:(\lambda,Y)\in(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\mapsto L_{\nu,t}^{\Lambda}(\lambda,Y)+I_{\nu}(Y).$ (2.17)
We also let
$K_{\nu,t}:(\Lambda,\lambda,Y)\in(-\infty,\ell_{\nu}]\times(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\mapsto J_{\nu,t}^{\Lambda}(\lambda,Y)-C_{t}-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}\,.$ (2.18)
@@
Let $\Lambda\leq\ell_{\nu}$. For every $\lambda<\ell_{\nu,t}$, we have
@@
The properties of $F^{+}_{\nu,t}$ listed at the end of the lemma are proven in Lemma B.2 and Lemma B.1, and we concentrate below on the proof of the functional inequality (2.19). Let $\varepsilon>0$. Let us perform the change of variables $X_{N}\mapsto(\lambda_{1},\ldots,\lambda_{N},v_{1},\ldots,v_{N})$ where $\lambda_{1}<\ldots<\lambda_{N}$ are the eigenvalues of $X_{N}$ and $v_{1},\ldots,v_{N}$ the associated eigenvectors. One has
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)=\frac{1}{Z_{N}^{t}}\int_{|\lambda_{1}-x|\leq\varepsilon}\prod_{i<j}|\lambda_{i}-\lambda_{j}|1_{\lambda_{1}<\ldots<\lambda_{N}}e^{-\frac{N}{4t}\mathrm{Tr}((OD(\lambda)O^{T}-B_{N})^{2})}\mathrm{d}O\mathrm{d}\lambda_{1}\ldots\mathrm{d}\lambda_{N},$
where $\mathrm{d}O$ is the Haar measure on $O_{N}(\mathbb{R})$, $D(\lambda)$ the diagonal matrix with entries $(\lambda_{i})_{1\leq i\leq N}$ and the constant $Z_{N}^{t}$ the value of the integral when the constraint on $\lambda_{1}$ is removed. Applying Lemma 2.2 to $M=B_{N}$ and using (2.9) gives, by the definition (2.16) of $L_{\nu,t}^{\Lambda}$,
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)=\frac{1}{Z^{t}_{N}}\int_{S^{N-1}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\\ \int_{|\lambda_{1}-x|\leq\varepsilon\atop\lambda_{1}<\ldots<\lambda_{N}}\prod_{i<j}|\lambda_{i}-\lambda_{j}|e^{-NL_{\nu,t}^{\Lambda}(\lambda_{1},\langle v_{1},w_{0}\rangle^{2},\ldots,\|P_{E_{\eta_{p}}}v_{1}\|^{2})-\frac{N}{4t}\lambda_{1}^{2}-\frac{N}{4t}\mathrm{Tr}(p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B_{N})p_{v_{1}}^{\perp})^{2}}\mathrm{d}\lambda_{1}\ldots\mathrm{d}\lambda_{N},$
where $\mathbb{P}_{v_{1}}$ is the Haar measure on the orthocomplement of $v_{1}$ and $D(0,\lambda)$ is the diagonal matrix with entries $(0,\lambda_{2},\ldots,\lambda_{N})$. If $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$, $d(\nu_{\alpha},\nu_{\mathsf{M}/N})\leq\delta$ and therefore $d(\nu_{\alpha}\boxplus\sigma_{t},\nu_{\mathsf{M}/N}\boxplus\sigma_{t})\leq\delta$. Hence, taking $\zeta=2\delta$ in (2.4), we have, uniformly on $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$,
@@
For every $\kappa>0$, set
@@
$\prod_{j:j>1}|\lambda_{1}-\lambda_{j}|\leq\exp\Bigr{\{}\sum_{i=1}^{N}f_{\kappa}(\lambda_{1}-\lambda_{i})\Bigr{\}}\leq\exp\Bigr{\{}N\int f_{\kappa}(\lambda_{1}-x)\mathrm{d}(\sigma_{t}\boxplus\nu)(x)+NC_{\kappa}\delta\Bigr{\}},$
@@
$\int f_{\kappa}(\lambda_{1}-x)\mathrm{d}(\sigma_{t}\boxplus\nu)(x)=\int\log|\lambda_{1}-x|\mathrm{d}(\sigma_{t}\boxplus\nu)(x)+O(\varepsilon)+O(\kappa)\,.$
Let $\varepsilon^{\prime}>0$. Since $I_{\nu}$ is lower semicontinuous, one can cover $\Delta^{p+1}$ by a finite union of measurable sets $\cup_{1\leq k\leq M}E_{Y_{k},\kappa_{k},\varepsilon^{\prime}}$ where for $Y\in\Delta^{p+1}$, and $\kappa,\varepsilon^{\prime}>0$,
[…]
and $(Y_{k})_{1\leq k\leq M}$ is a family of points of $\Delta^{p+1}$. We denote in short $E_{Y_{k}}$ for $E_{Y_{k},\kappa_{k},\varepsilon^{\prime}}$. We deduce by continuity of $L_{\nu,\Lambda}$ together with (2.20), (2.21) that
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)\leq\frac{2\varepsilon}{Z_{N}^{t}}\sum_{k=1}^{M}e^{N(\int\log|x-y|\mathrm{d}(\sigma_{t}\boxplus\nu)(x)-L_{\nu,t}^{\Lambda}(x,Y_{k})+o_{\kappa}(\delta,\epsilon,\epsilon^{\prime})-\frac{1}{4t}x^{2})}\int_{E_{Y_{k}}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\qquad\qquad$
$\qquad\int_{\mathcal{B}(\delta,\varepsilon,x)}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}+O(e^{-\min\{C_{\delta},1\}N^{2}})\,,$ (2.23)
where $o_{\kappa}(\delta,\epsilon,\epsilon^{\prime})=O(\varepsilon)+O(\varepsilon^{\prime})+C_{\kappa}(\delta)+O(\kappa)$ and
$\mathcal{B}(\delta,\varepsilon,x):=\{\lambda_{1}<\cdots<\lambda_{N}\}\cap\{|\lambda_{1}-x|\leq\varepsilon\}\cap\Bigr{\{}d\Bigr{(}\frac{1}{N}\sum\delta_{\lambda_{i}},\nu\boxplus\sigma_{t}\Bigr{)}\leq\delta\Bigr{\}}\cap\Bigr{\{}\frac{1}{N}\sum\lambda_{i}^{2}\leq L_{0}+1\Bigr{\}}\,.$
Fix $Y\in\Delta^{p+1}$ and $v_{1}\in E_{Y,\kappa,\varepsilon}$. We can write
$\int_{\mathcal{B}(\delta,\varepsilon,x)}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\cdots\mathrm{d}\lambda_{N}\\ \leq\int_{x-\varepsilon\leq\lambda_{2}\leq\cdots\leq\lambda_{N}}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B_{N})p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\cdots\mathrm{d}\lambda_{N}.$
Let us remark that $OD(0,\lambda)O^{T}$ follows the law of a GOE of dimension $N-1$ and variance $t(N-1)/N$ so that
$\int\mathrm{d}\mathbb{P}_{v_{1}}(v_{2},\ldots,v_{N})\int_{\lambda_{2}\leq\cdots\leq\lambda_{N}}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B)p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}=Z_{N-1}^{t\frac{N-1}{N}}\,.$
@@
from which the limit of $\frac{1}{N}\log\frac{Z_{N-1}^{t\frac{N-1}{N}}}{Z_{N}^{t}}$ is easily inferred and gives a contribution $C_{t}$ with $C_{t}$ as in (2.15).
@@
By Lemma 2.4, the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ has eigenvalues $\eta_{1},\ldots,\eta_{p}$ of multiplicities between $M_{k}-1$ and $M_{k}$ with $p$ possible outliers $\Lambda_{1}\leq\cdots\leq\Lambda_{p}$ with for each $k=2,\ldots,p$, $\Lambda_{k}\in[\eta_{k-1},\eta_{k}]$. Moreover, by the continuity of $\Lambda_{1}$ proven in Lemma 2.4, $\Lambda_{1}$ is at distance at most of order $\varepsilon$ from $\Phi(\Lambda,Y)$ if $v_{1}\in E_{Y,\kappa,\varepsilon^{\prime}}$ and $\kappa$ is small enough. We can bound $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ from above by a matrix $B_{N}^{\prime}$ in $v_{1}^{\perp}$ with the same eigenspaces, the outlier $\Lambda_{1}$ and eigenvalues $\eta_{i}$ with multiplicity $M_{i}^{\prime}=M_{i}$ except for $\eta_{1}$ which has multiplicity $M_{1}^{\prime}=M_{1}-1$. We therefore get
@@
$\frac{1}{Z_{N-1}^{t\frac{N-1}{N}}}\int_{x-\varepsilon\leq\lambda_{2}\leq\cdots\leq\lambda_{N}}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((p_{v_{1}}^{\perp}(OD(0,\lambda)O^{T}-B_{N})p_{v_{1}}^{\perp})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}\qquad$
$\qquad\leq\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N}^{\prime})\geq x-2\varepsilon),$
@@
where we incorporated the error due to the small change in the variance (which can be removed after rescaling and boundedness of $B_{N}$) in an additional error of order $\varepsilon$ for $N$ large enough. Note that Theorem 1.1 and Lemma 2.4 imply that the smallest eigenvalue of $G_{N-1}+B_{N}^{\prime}$ converges towards $H_{\nu,t}(\Lambda_{1})\simeq H_{\nu,t}(\Phi(\Lambda,Y))$ if $\Phi(\Lambda,Y)<\omega_{\nu,t}(\ell_{\nu,t})$ and towards $\ell_{\nu,t}$ otherwise. We therefore have three cases:

### v2

@@ 2.3. Outliers of $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$
$\frac{y_{0}(v_{1})}{\Lambda_{1}-\Lambda}+\sum_{k=1}^{p}\frac{y_{k}(v_{1})}{\Lambda_{1}-\eta_{k}}=0.$
$\Lambda_{1}$ is thus the unique solution of (2.15) as claimed. Item (3) is clear. ∎
We next study the typical behavior of the smallest eigenvalue of $p_{v_{1}}^{\perp}(G_{N}+B_{N})p_{v_{1}}^{\perp}$ in $v_{1}^{\perp}$, when $v_{1}$ is as in Lemma 2.4. It is a straightforward application of Lemma 2.4 and Theorem 1.1 as this matrix is the sum of a GOE matrix and the matrix $B_{N-1}^{\prime}$ with spectrum asymptotically distributed according to $\nu$ and outlier $\Phi(Y,\Lambda)$ as in Lemma 2.4.
@@ 2.4. A functional inequality for the large deviation upper bound
Let $B_{N}$ be a diagonal matrix with eigenvalues $\Lambda$ with multiplicity $1$, $\eta_{1},\ldots,\eta_{p}$ of multiplicities $M_{1},\ldots,M_{p}$. Let $\mathsf{M}=(M_{1},\ldots,M_{p})\in\mathcal{B}_{\delta}^{N}(\alpha)$. For every $\Lambda\leq\ell_{\nu}$, let us denote $\mathbb{P}_{N,\Lambda,\mathsf{M}}$ the law of $X_{N}=G_{N}+B_{N}$ and
@@
$F_{\nu,t}^{+}(\Lambda,\lambda):=-\limsup_{\Lambda^{\prime}\to\Lambda}\lim_{\varepsilon\downarrow 0}\lim_{\delta\downarrow 0}\limsup_{N\to\infty}\sup_{\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)}\frac{1}{N}\log P_{N,\Lambda^{\prime},\mathsf{M}}(\lambda,\varepsilon).$ (2.18)
@@
$g:x\in\mathbb{R}\mapsto\limsup_{y\rightarrow x}f(y)=\lim_{\varepsilon\to 0}\sup_{B(x,\varepsilon)}f$
@@
Recall the map $L_{\nu,t}^{\Lambda}$ defined in (2.9) and the map $I_{\nu}:Y\in\Delta^{p+1}\mapsto-\frac{1}{2}\sum_{k=1}^{p}\alpha_{k}\log(y_{k}/\alpha_{k})$ from Lemma 2.3. We set
$J_{\nu,t}^{\Lambda}:(\lambda,Y)\in(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\mapsto L_{\nu,t}^{\Lambda}(\lambda,Y)+I_{\nu}(Y),$ (2.20)
and, for $(\Lambda,\lambda,Y)\in(-\infty,\ell_{\nu}]\times(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}$, we define
$K_{\nu,t}(\Lambda,\lambda,Y)=J_{\nu,t}^{\Lambda}(\lambda,Y)-C_{t}-\int\log|\lambda-x|\mathrm{d}(\nu\boxplus\sigma_{t})(x)+\frac{\lambda^{2}}{4t}\,.$ (2.21)
@@
Let $\Lambda\leq\ell_{\nu}$. Let $\Phi(\Lambda,\cdot)$ be as in (2.14) and $\ell_{\nu,t}^{\Lambda}$ be as in (1.14). For every $\lambda<\ell_{\nu,t}$, we have
@@
The properties of $F^{+}_{\nu,t}$ listed at the end of the lemma are proven in Lemma B.2 and Lemma B.1, and we concentrate below on the proof of the functional inequality (2.22). If $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$, $d(\nu_{\alpha},\nu_{\mathsf{M}/N})\leq\delta$ and therefore $d(\nu_{\alpha}\boxplus\sigma_{t},\nu_{\mathsf{M}/N}\boxplus\sigma_{t})\leq\delta$. Hence, taking $\zeta=2\delta$ in (2.4), we have, uniformly on $\mathsf{M}\in\mathcal{B}_{\delta}^{N}(\alpha)$,
@@
Therefore, if for $L>L_{0}$ and $\delta>0$ fixed, we let
$K_{L,\delta}=\{\lambda^{N}\in\mathbb{R}^{N}:\frac{1}{N}\sum\lambda_{i}^{2}\leq L,d\Bigr{(}\frac{1}{N}\sum\delta_{\lambda_{i}},\nu_{\alpha}\boxplus\sigma_{t}\Bigr{)}\leq\delta\},$
then we deduce from (2.24) and (2.23) that
$P_{N,\Lambda,\mathsf{M}}(x,\varepsilon)=\mathbb{P}_{N,\Lambda,\mathsf{M}}\Bigr{(}|\lambda_{1}(X_{N})-x|\leq\varepsilon,\lambda^{N}(X_{N})\in K_{L,\delta})+O(e^{-c(L,\delta)N^{2}})$ (2.25)
for some positive real number $c(\delta,L)$. We denote in the following
$\mathcal{C}_{L,\delta}(x,\varepsilon):=\{\lambda^{N}\in K_{L,\delta}:\lambda_{1}<\lambda_{2}\cdots<\lambda_{N},|\lambda_{1}-x|\leq\varepsilon\}\,.$
By (2.25), one needs to upper bound
$P^{L,\delta}_{N,\Lambda,\mathsf{M}}(x,\varepsilon):=\mathbb{P}_{N,\Lambda,\mathsf{M}}\Bigr{(}\lambda^{N}(X_{N})\in\mathcal{C}_{L,\delta}(x,\varepsilon)\Bigr{)}\,.$
Let us perform the change of variables $X_{N}\mapsto(\lambda_{1},\ldots,\lambda_{N},v_{1},\ldots,v_{N})$ where $\lambda_{1}<\ldots<\lambda_{N}$ are the eigenvalues of $X_{N}$ and $v_{1},\ldots,v_{N}$ the associated eigenvectors. We find
$P^{L,\delta}_{N,\Lambda,\mathsf{M}}(x,\varepsilon)=\frac{1}{Z_{N}^{t}}\int_{\lambda^{N}\in\mathcal{C}_{L,\delta}(x,\varepsilon)}\prod_{i<j}|\lambda_{i}-\lambda_{j}|1_{\lambda_{1}<\ldots<\lambda_{N}}e^{-\frac{N}{4t}\mathrm{Tr}((O_{N}D_{N}(\lambda)O_{N}^{T}-B_{N})^{2})}\mathrm{d}O_{N}\mathrm{d}\lambda^{N},$
where $\mathrm{d}O_{N}$ is the Haar measure on $O_{N}(\mathbb{R})$, $D_{N}(\lambda)$ the diagonal matrix with entries $(\lambda_{i})_{1\leq i\leq N}$ and the constant $Z_{N}^{t}$ the value of the integral when the constraint on $\lambda_{1}$ is removed. $\mathrm{d}\lambda^{N}$ denotes $\mathrm{d}\lambda_{1}\ldots\mathrm{d}\lambda_{N}$ in short. Applying Lemma 2.2 to $M=B_{N}$ and using (2.12) gives, by the definition (2.9) of $L_{\nu,t}^{\Lambda}$,
$P_{N,\Lambda,\mathsf{M}}^{L,\delta}(x,\varepsilon)$ $=$ $\int_{S^{N-1}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(O_{N-1})\mathbb{M}(\mathcal{C}_{L,\delta}(x,\varepsilon),v_{1},O_{N-1})$
$\mathbb{M}(A,v,O)$ $=$ $\ \frac{1}{Z^{t}_{N}}\int_{\lambda^{N}\in A}\prod_{i<j}|\lambda_{i}-\lambda_{j}|e^{-NL_{\nu,t}^{\Lambda}(\lambda_{1},Y(v))-\frac{N}{4t}\lambda_{1}^{2}-\frac{N}{4t}\mathrm{Tr}(OD_{N-1}(\lambda)O^{T}-B_{N-1}^{\prime})^{2}}\mathrm{d}\lambda^{N}$
where $\mathbb{P}_{v_{1}}$ is the Haar measure on the orthocomplement of $v_{1}$, $B_{N-1}^{\prime}$ the projection of $B_{N}$ onto the orthocomplement of $v_{1}$ and $D_{N-1}(\lambda)$ is the diagonal matrix with entries $(\lambda_{2},\ldots,\lambda_{N})$.
Let $\varepsilon^{\prime},\kappa,\zeta>0$. We denote $E_{Y_{0}}=\{y\in\Delta^{p+1}:y_{1}<2\zeta\}$. Since $I_{\nu}$ is lower semicontinuous, one can cover $E_{Y_{0}}^{c}$ by a finite union of measurable sets $\cup_{1\leq k\leq M}E_{Y_{k},\kappa_{k},\varepsilon^{\prime}}$ where for $Y\in\Delta^{p+1}$, and $\kappa,\varepsilon^{\prime}>0$,
[…]
and $(Y_{k})_{1\leq k\leq M}$ is a family of points of $\Delta^{p+1}$. We can assume from the start that $\kappa_{j}\leq\zeta$ for all $j$ so that $Y_{j}(1)\geq\zeta>0$. We denote in short $E_{Y_{k}}$ for $E_{Y_{k},\kappa_{k},\varepsilon^{\prime}}$. We therefore find
$P_{N,\Lambda,\mathsf{M}}^{L,\delta}(x,\varepsilon)=\sum_{k=0}^{M}\int_{v_{1}\in E_{Y_{k}}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(O_{N-1})\mathbb{M}(\mathcal{C}_{L,\delta}(x,\varepsilon),v_{1},O_{N-1})$ (2.27)
We next estimate each term in the above right hand side. For every $\kappa>0$, set
@@
$\prod_{j:j>1}|\lambda_{1}-\lambda_{j}|\leq\exp\Bigr{\{}\sum_{i=1}^{N}f_{\kappa}(\lambda_{1}-\lambda_{i})\Bigr{\}}\leq\exp\Bigr{\{}N\int f_{\kappa}(\lambda_{1}-y)\mathrm{d}(\sigma_{t}\boxplus\nu)(y)+NC_{\kappa}\delta\Bigr{\}},$
@@
$\int f_{\kappa}(\lambda_{1}-x)\mathrm{d}(\sigma_{t}\boxplus\nu)(x)=\int\log|x-y|\mathrm{d}(\sigma_{t}\boxplus\nu)(y)+O(\varepsilon)+O(\kappa)\,.$
Similarly $\lambda_{1}^{2}/4t$ is very close to $x^{2}/4t$. As a consequence, we find that
$\mathbb{M}(\mathcal{C}_{L,\delta}(x,\varepsilon),v_{1},O_{N-1})=e^{N(\int\log|x-y|\mathrm{d}(\sigma_{t}\boxplus\nu)(y)-\frac{x^{2}}{4t}+o(1))}\mathbb{L}(\mathcal{C}_{L,\delta}(x,\varepsilon),v_{1},O_{N-1})$ (2.28)
with
$\mathbb{L}(A,v_{1},O_{N-1})=\frac{1}{Z^{t}_{N}}\int_{\lambda^{N}\in A}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-NL_{\nu,t}^{\Lambda}(\lambda_{1},Y(v_{1}))-\frac{N}{4t}\mathrm{Tr}(O_{N-1}D_{N-1}(\lambda)O_{N-1}^{T}-B_{N-1}^{\prime})^{2}}\mathrm{d}\lambda^{N}\,.$
Let us finally remark that $O_{N-1}D_{N-1}(\lambda)O_{N-1}^{T}$ follows the law of a GOE of dimension $N-1$ and variance $t(N-1)/N$ so that
$\int\mathrm{d}\mathbb{P}_{v_{1}}(O_{N-1})\int_{\lambda_{2}\leq\cdots\leq\lambda_{N}}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((O_{N-1}D_{N-1}(\lambda)O_{N-1}^{T}-B_{N-1}^{\prime})^{2})}\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}=Z_{N-1}^{t\frac{N-1}{N}}\,.$ (2.29)
@@
from which the limit of $\frac{1}{N}\log\frac{Z_{N-1}^{t\frac{N-1}{N}}}{Z_{N}^{t}}$ is easily computed to be equal to $C_{t}$ as in (2.19).
We first estimate the first term in the right hand side of (2.27). Putting together (2.28) and (2.29) we see that there is a finite constant $C$ so that
$\int_{v_{1}\in E_{Y_{0}}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(O_{N-1})\mathbb{M}(\mathcal{C}_{L,\delta}(x,\varepsilon),v_{1},O_{N-1})\leq e^{CN}\mathbb{P}(y_{1}(v_{1})\leq 2\zeta)\leq e^{-C^{\prime}N}(\frac{2\zeta}{\alpha_{1}})^{\frac{1}{2}\alpha_{1}N}$
where we finally used Lemma 2.3 and assumed $N$ enough.
@@
We now estimate the next terms in the right hand side of (2.27). When $v_{1}\in E_{Y,\kappa,\varepsilon^{\prime}}$ for $k\geq 1,$ $Y\in E_{Y_{0}}^{c}$ and $|\lambda_{1}-x|\leq\varepsilon$, $L_{\nu,t}^{\Lambda}(\lambda_{1},Y(v_{1}))$ is at distance from $L_{\nu,t}^{\Lambda}(x,Y_{k})$ of order $\kappa+\varepsilon$. Moreover, $y_{1}(v_{1})\geq\zeta$ if $\kappa\leq\zeta$. By Lemma 2.4, the matrix $B_{N-1}^{\prime}$ has eigenvalues $\eta_{1},\ldots,\eta_{p}$ of multiplicities between $M_{k}-1$ and $M_{k}$ with $p$ possible outliers $\Lambda_{1}\leq\cdots\leq\Lambda_{p}$ with for every $k=2,\ldots,p$, $\Lambda_{k}\in[\eta_{k-1},\eta_{k}]$. Moreover, by the continuity of $\Lambda_{1}$ proven in Lemma 2.4, $\Lambda_{1}$ is at distance at most of order $\kappa$ from $\Phi(\Lambda,Y)$. We can bound $B_{N-1}^{\prime}$ from above by a matrix $B_{N-1}^{\prime\prime}$ in $v_{1}^{\perp}$ with the same eigenspaces, the outlier $\Lambda_{1}$ and eigenvalues $\eta_{i}$ with multiplicity $M_{i}^{\prime}=M_{i}$ except for $\eta_{1}$ which has multiplicity $M_{1}^{\prime}=M_{1}-1$. We therefore get
$\int_{v_{1}\in E_{Y_{k}}}\mathrm{d}v_{1}\int\mathrm{d}\mathbb{P}_{v_{1}}(O_{N-1})\mathbb{M}(\mathcal{C}_{L,\delta}(x,\varepsilon),v_{1},O_{N-1})\qquad\qquad\qquad\qquad$
[…]
where $\mathbb{K}(x,\varepsilon,v_{1})$ is equal to
@@
$=$ $\frac{1}{Z_{N-1}^{t\frac{N-1}{N}}}\int\mathrm{d}\mathbb{P}_{v_{1}}(O_{N-1})\int_{\tiny x-2\varepsilon\leq\lambda_{2}\leq\cdots\leq\lambda_{N}}\prod_{2\leq i<j}|\lambda_{i}-\lambda_{j}|e^{-\frac{N}{4t}\mathrm{Tr}((O_{N-1}D_{N-1}(\lambda)O_{N-1}^{T}-B_{N-1}^{\prime})^{2})}\mathrm{d}\lambda^{N-1}$
$=$ $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(\frac{N}{N-1}G_{N-1}+B_{N-1}^{\prime})\geq x-2\varepsilon)$
$\leq$ $\mathbb{P}_{N-1,\Lambda_{1},\mathsf{M}^{\prime}}(\lambda_{1}(G_{N-1}+B_{N-1}^{\prime\prime})\geq x-3\varepsilon)=:\mathbb{K}(x,\varepsilon,Y_{k})$
@@
where we denoted $\mathrm{d}\lambda^{N-1}=\mathrm{d}\lambda_{2}\ldots\mathrm{d}\lambda_{N}$, and in the last line we incorporated the error due to the small change in the variance (which can be removed after rescaling and boundedness of $B_{N}^{\prime}$) in an additional error of order $\varepsilon$ for $N$ large enough. Note here that $\mathbb{K}(x,\varepsilon,Y_{k})$ only depends on $Y_{k}$.
Note that Theorem 1.1 and Lemma 2.4 imply that, if $v_{1}$ belongs to $E_{Y,\kappa,\epsilon}$ and the empirical measure of the $\lambda_{i}$ is at distance smaller than $\delta$ from $\nu\boxplus\sigma_{t}$, the smallest eigenvalue of $G_{N-1}+B_{N-1}^{\prime\prime}$ is close to $H_{\nu,t}(\Lambda_{1})\simeq H_{\nu,t}(\Phi(\Lambda,Y))$ if $\Phi(\Lambda,Y)<\omega_{\nu,t}(\ell_{\nu,t})$ and close to $\ell_{\nu,t}$ otherwise, with some error going to zero with $\delta$ and $\kappa$. We therefore have three cases: